OpenStudy (anonymous):

If A^(x) = 3, find A^(10x-x) - 5

OpenStudy (lgbasallote):

\[\huge A^{10x - x} \implies \frac{A^{10x}}{A^x}\]

OpenStudy (lgbasallote):

\[\huge A^{10x - x} - 5 \implies \frac{A^{10x}}{A^x} - 5\]

OpenStudy (lgbasallote):

\[\huge A^{10x} \implies (A^x)^{10}\]

OpenStudy (lgbasallote):

\[\huge \frac{A^{10x}}{A^x} - 5 \implies \frac{(A^x)^{10}}{A^x} - 5\] does that help?

OpenStudy (lgbasallote):

did you get my point @deehthebee ??

OpenStudy (anonymous):

@deehthebee answer him he is asking you something.. Interact with him..

OpenStudy (anonymous):

Sorry. First time!

OpenStudy (anonymous):

Yes, that was very clear. Thank you so much

OpenStudy (lgbasallote):

so can you do the rest? :)

OpenStudy (anonymous):

I'll try to. I'll come back if I need any help

OpenStudy (lgbasallote):

sure ^_^

OpenStudy (anonymous):

Sidenote: @lgbasallote Why didn't you just simplify in the beginning (10x-x=9x)?

OpenStudy (lgbasallote):

because the question didnt simplify it :(

OpenStudy (lgbasallote):

i assume they wanna be fancy so i did

OpenStudy (lgbasallote):

i think this is an exponent rule application...so i tried applying

OpenStudy (anonymous):

So would it be (A^x)^9 - 5?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

Thank you so, so much!

OpenStudy (lgbasallote):

\[\huge \color{maroon}{\mathtt{\text{<tips hat>}}}\]

OpenStudy (anonymous):

Sorry I'm back. If A^x = 3, wouldn't the answer be 19678?

OpenStudy (lgbasallote):

wait...ill use me calculator

OpenStudy (lgbasallote):

correct!

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Can someone explain the mathematics behind doing some of these steps though?

OpenStudy (lgbasallote):

which ones?

OpenStudy (anonymous):

Just the first one

OpenStudy (lgbasallote):

\[A^{10x-x} \implies \frac{A^{10x}}{A^x}\] that one?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Sorry for the delay, the message didn't get through at first

OpenStudy (anonymous):

\[\LARGE{(x^{3})(x^{4}) = (x*x*x)*(x*x*x*x) = x^{7}}\] You add exponents when multiplying similar terms. The other way around also works: \[\LARGE{x^{7} = (x^{3})(x^{4})}\] \[\LARGE{x^{7} = x^{6+1} = x^{6}x^{1}}\] \[\LARGE{x^{7} = x^{8-1} = x^{8}x^{-1}}\] Negative exponents are another way of saying that you're multiplying by the inverse, so: \[\LARGE{x^{8}x^{-1} = x^{8}\frac{1}{x^{1}} = \frac{x^{8}}{x^{1}} = \frac{x^{7+1}}{x^{1}}= \frac{x^{7}x^{1}}{x^{1}} = x^{7}}\] It's just quicker and easier to say 10x-x=9x though :)

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

And knowing is half the battle, GI JOE!!!!!!!!!!

OpenStudy (lgbasallote):

haha nice..that was pretty long @Wired

OpenStudy (anonymous):

TY